Given that one of the roots of the quadratic equation 4x^2-(p-2)x-2p=5 is negative of the other root,find
(a)the value of p,
(b)the roots of the quadratic equation

Answer by stanbon(75887)

Given that one of the roots of the quadratic equation 4x^2-(p-2)x-2p=5 is negative of the other root,find
(a)the value of p,
Let x and -x be the roots.
Then,
4x^2-(p-2)x-2p-5 = 0
And, replacing x with -x,
4x^2 + (p-2)x -2p -5 = 0
-----------------------------------
Subtracting the 1st from the 2nd you get:
2(p-2)x = 0
---
So x = 0 or p = 2
===========================
(b)the roots of the quadratic equation
4x^2-(p-2)x-2p=5
If p =2
4x^2 -2p = 5
4x^2= 2p+5
x^2 = (2p+5)/4
x = +-sqrt[(2p+5)/4]
Since p = 2
x = +-sqrt[9/4]
x = +-3/2
=================================
Cheers,
Stan H.
--------------------------